OpenStudy (anonymous):

What is the solution set for |–4x – 8| = 4?

OpenStudy (anonymous):

\[|-4x-8|=4\] \[-4x-8=4\] or \[-4x-8=-4\] first one gives \[-4x-8=4\] \[-4x=12\] \[x=-3\]second one gives \[-4x-8=-4\] \[-4x=4\] \[x=-1\]